Now suppose a reaction vessel is filed with 1.98 atm of dintrogen tetroxide (N-04) at 341. C. Answer the following questions about this system fall ts it possible to reverse this tendency by adding NO2 In other words, If you said the pressure of N 0, will tend to rise, can that be changed to a tendency to fall by adding NO2? Similarty, it you said the pressure of N204 will tend to fall, can that be changed to a tendency to rise by adding NO. the minimum pressure of NO2 needed to reverse it 囗
Answer #1

We start with pure N_2O_4. Some of it will dissociate into N O_2 to reach equilibrium.

Hence, the pressure of N_2O_4 will fall. If we add some N O_2, it will associate to form N_2O_4.

This will increase the pressure of N_2O_4. This will reverse the tendency.

\\ \: \\ \: \Delta G^0 = -5.4 \: kJ = -5400 \: J \\ \: \\ \: \Delta G^0 = -RTlnK \\ \: \\ \: -5400 \: J = -8.314 \: J/mol/K \times (341.+273) \: K \times lnK \\ \: \\ \: lnK=1.058 \\ \: \\ \:K=2.88

\\ \: \\ \:K= \dfrac { P_{N2O4} }{ P_{NO2}^2 } \\ \: \\ \: 2.88=\dfrac { 1.98 }{ P_{NO2}^2 } \\ \: \\ \:P_{NO2}^2 = 0.688 \\ \: \\ \:P_{NO2}=0.829

When the pressure of NO2 exceeds, the tendency can be reversed.

Consider the following equilibrium: Now suppose a reaction vessel is filled with 1.98 atm of dinitrogen tetroxide (N204) at 3